1)Find the slope of the line containing
P1 (12, −8) and P2 (10, −14).

2)Find the slope of the line containing the points P1(5, 5) and P2(9, 2).

Answer by Maths68(1474)

1)Find the slope of the line containing
P1 (12, −8) and P2 (10, −14).
Solution
Slope of line m = y2-y1/x2-x1
(x1,y1)=(12, −8) and (x2,y2)= (10, −14).
m = (-14-(-8)) / (10-12)
m=(-14+8)/-2
m=-6/-2
m=3
Slope m = 3

2)Find the slope of the line containing the points P1(5, 5) and P2(9, 2).
(x1,y1)=(5, 5) and (x2,y2)= (9, 2)
Slope of line m = y2-y1/x2-x1
Slope of line m = (2-5)/(9-5)
Slope of line m = -3/4